OpenStudy (anonymous):

Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 19 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (jdoe0001):

if you were to solve the 1st equation for "x", what would that look like?

OpenStudy (amistre64):

accursed -y lol let x=8+n let y=-8+n and solve for n

OpenStudy (anonymous):

What?

OpenStudy (amistre64):

your going to have to be more descriptive about your confusion ....

OpenStudy (anonymous):

y= -12 and x = 7 because of substitution: y = 2-2x so x- (2-2x)= 19 x-2+2x=19 3x-2=19 3x=21 x=7 and y = 2-2*7=-12